OpenStudy (anonymous):

Calculate the gradient of the scalar function ∅=e^(xy^2z^3)

OpenStudy (anonymous):

df/dx = (y^2z^3)e^(xy^2z^3) df/dy = (2xyz^3)e^(xy^2z^3) df/dz = (3xy^2z^2)e^(xy^2z^3) so grad f = ( (y^2z^3)e^(xy^2z^3) , (2xyz^3)e^(xy^2z^3) , (3xy^2z^2)e^(xy^2z^3))

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Me: "Mua ha ha! I'm in Cal 2! Nothing can stop me!" *visits open study after a 1 year hiatus and sees this problem* Me: "Nope." :P